OpenStudy (anonymous):

If the revenue of a cookie company is f(x) = 2.5x + 16, where x is the number of cookies sold, and the expenses of the company are g(x) = x + 20, where x is also the number of cookies sold, what is the profit of the company if they sell 126 cookies?

OpenStudy (perl):

profit = revenue - expenses. so p(x) = f(x) - g(x) find p(126) p(126) = f(126) - g(126)

OpenStudy (anonymous):

so that means the profit 185.00